OpenStudy (anonymous):

The total cost to park a car in a parking garage that has an entry fee of $4 can be determined by using the equation c = 2h + 4 where c is the total cost and h is the number of hours the car is in the garage. The garage informed Robert that the entry fee had changed and that his bill for parking his car for 6 hours is $20. Which equation models the cost to park his car for h hours with the new entry fee?

OpenStudy (anonymous):

A. c = 2h + 20 B. c = h + 6 C. c = 2h + 8 D. c = h + 14